OpenStudy (anonymous):

Find the slope of the line that passes through (3, -4) and (-1, 2)

OpenStudy (earthcitizen):

m = -1.5

OpenStudy (anonymous):

that not right but thank you any way

OpenStudy (anonymous):

1 points Save Find the slope of the line passing through (4, 9) and (-2, 9).

OpenStudy (earthcitizen):

\[m = (y _{2} - y _{1})/(x _{2}-x _{1}) = (2+4)/(-1-3) = -3/2 = -1.5\]

OpenStudy (anonymous):

this is not right because it not on my answer sheet so idk what to do.

OpenStudy (anonymous):

rise over run, so when you draw the graph and count the amount that it rises and divide it by the amount that it runs, then you get [6\div-4\] (6/-4) which is -1.5

OpenStudy (anonymous):

Then your answer sheet is wrong honestly, unless you gave us the wrong points.

OpenStudy (earthcitizen):

for the second question m = 0

OpenStudy (anonymous):

i got that one thank you i need help on other one."

OpenStudy (anonymous):

Write the direct variation of the equation if y varies directly with x and y = 45 when x = 9.

OpenStudy (earthcitizen):

\[y \alpha x => y = kx\] k =y/x

OpenStudy (anonymous):

thank you s much

OpenStudy (anonymous):

you misspelled so.

OpenStudy (earthcitizen):

no p

OpenStudy (anonymous):

Which of the following represents the solution of x - 8 < -2 in set notation? Question 3 answers